# A Knowledge Graph-based Retrieval-Augmented Generation Framework for Algorithm Selection in the Facility Layout Problem

### v1

@@ Abstract
Selecting a solution algorithm for the Facility Layout Problem (FLP), an NP-hard optimization problem with a multiobjective trade-off, is a complex task that requires deep expert knowledge. The performance of a given algorithm depends on specific problem characteristics such as its scale, objectives, and constraints. This creates a need for a data-driven recommendation method to guide algorithm selection in automated design systems. This paper introduces a new recommendation method to make such expertise accessible, based on a Knowledge Graph-based Retrieval-Augmented Generation (KG-RAG) framework. To address this, a domain-specific knowledge graph is constructed from published literature. The method then employs a multi-faceted retrieval mechanism to gather relevant evidence from this knowledge graph using three distinct approaches, which include a precise graph-based search, flexible vector-based search, and high-level cluster-based search. The retrieved evidence is utilized by a Large Language Model (LLM) to generate algorithm recommendations with data-driven reasoning. The proposed KG-RAG method is compared against a commercial LLM chatbot with access to the knowledge base as a table, across a series of diverse, real-world FLP test cases. Based on recommendation accuracy and reasoning capability, the proposed method performed significantly better than the commercial LLM chatbot.
Keywords: Facility Layout Problem, Algorithm Selection, Knowledge Graph, Retrieval-Augmented Generation (RAG), Large Language Models (LLM), Recommender Systems.
@@ 1 Introduction
The Facility Layout Problem (FLP) is an important part of industrial engineering and operations research. [1] Because the operating cost and service performance are directly influenced by effective layout design. The suitable placement of facilities reduces the total operating costs by 20% to 50% [2]. The computational complexity of FLP stems from its combined discrete and continuous nature. In practice, this complexity is magnified by multi-objective trade-offs. For example, minimize material handling costs and maximize safety. This multi-objective problem is coupled with constraints such as aspect-ratio, adjacency and circulation constraints, and dynamic considerations such as reconfigurations to accommodate product-mix changes, demand changes, or equipment changes. These considerations render many FLP variants NP-hard, so exact mathematical programming is computationally tractable only for small or specially structured instances [3]. This FLP is also dynamic in nature because of evolving real-world demand. The problem gets further complicated because selecting a suitable optimisation algorithm for this complex problem is challenging and requires expert knowledge.
@@
The Solution techniques have been evolving as the complexity of the FLP increases, as mentioned above. Initial solutions, such as quadratic assignment problem [4], mixed-integer programming and graph-theoretic formulations [5] for equal-area FLP. These are optimum guaranteed, but with exponential growth in computational time, limiting industrial applications to problems. As the problem size grows, researchers shifted to heuristic approaches, such as construction and improvement heuristics. Here, the construction heuristics build layouts from the ground up, for example, automated layout design program and computerized relationship layout planning [6] whereas the improvement heuristics build solutions incrementally, for example computerized relative allocation of facilities technique [6]. These techniques are likely to converge to suboptimal layouts. To enhance the ability of handling large and complex search spaces, various metaheuristic techniques, including particle swarm optimization, genetic algorithms, ant simulated annealing, colony optimization, and others, have been introduced [7]. These techniques increase the likelihood of avoiding local optima. Moreover, the hybrid metaheuristic approaches, such as biased random-key genetic algorithms with linear programming repair, have recently been developed as state-of-the-art solutions, with enhanced solution quality and scalability [8].
@@
Despite these developments, the situation in FLP today is a broad and scattered collection of niche algorithms. This is typical of the no free lunch theorems, which state that when performance is averaged across all possible optimization problems, no single algorithm is universally superior. The theorems imply that an algorithm performs better on one class of problems and may perform weaker on others. For example, an algorithm that performs well in small-scale, static settings may be desired in large-scale, multi-objective, or dynamic conditions [9]. As a result, attention has turned to the meta-problem of algorithm selection, which involves selecting the algorithm, representation, and parameters best suited to an instance of information [21].
@@
The early attempts to address this meta-problem were static decision-support methods. For instance, a GenOpt-based framework to guide practitioners in selecting from a wide range of general optimization algorithms. Their approach, which utilized flowcharts and choice matrices, was designed to assist novices by matching algorithm characteristics to problem types and performance requirements. While this represents a valuable step towards systematizing algorithm selection, the framework was not specific to the unique complexities of the FLP. Furthermore, like other static methods, it is not dynamic and cannot adapt to new algorithms or shifting problem definitions without manual updates [10]. To improve flexibility, research employed data-driven approaches, namely meta-learning, that predict algorithm performance from problem meta-features, for example, size, constraints, and statistical properties [11]. SATzilla is just an example, where it selects the fastest solver from a portfolio based on instance features and performs well in combinatorial optimisation [12]. This has evolved to AutoML, where algorithm selection is combined with hyperparameter tuning in the combined algorithm selection and hyperparameter framework [13]. While adaptive frameworks and AutoML have built significant capability, Large Language Models (LLMs) such as OpenAI’s GPT series and Google’s Gemini introduce new possibilities [14]. Large, general-purpose corpora-trained, these models are well-suited to acquiring language, reasoning, and in-context learning but fail on domain-specific problems such as FLP optimisation, hallucinations (providing plausible but incorrect responses) [15] and weak domain adaptation [16].
@@
AI research has turned to grounding LLMs with structured, verifiable knowledge to address these weaknesses. Two key paradigms dominate Retrieval-Augmented Generation (RAG) and Knowledge Graphs (KG). RAG retrieves relevant, domain-specific content before generation, reducing hallucinations and enabling integration of up-to-date information without retraining [17]. KG organise domain entities and relationships, enabling explainable reasoning and mitigating data sparsity or cold-start issues. The relational pathways in KG enhance accuracy and interpretability, which are critical for decision support in technical domains [18]. A new approach integrates these paradigms. A KG-RAG [19] model uses a structured KG as the retrieval source for a RAG. The KG offers an explainable, linked domain model, and RAG allows LLMs to flexibly query and blend this knowledge into transparent, evidence-based proposals. This integration is suitable for comprehensive domain knowledge and justifiable output for complex recommendation tasks. The literature indicates a convergence of a long-standing operations research problem with a contemporary AI opportunity. FLP algorithmic fragmentation under the No Free Lunch principle renders selection a vital bottleneck. Current machine learning-based selectors, though adaptive, tend to operate as black boxes. Concurrently, the KG-RAG paradigm addresses LLM limitations of hallucinations and poor performance in domain adaptation while allowing for structured, transparent reasoning. KG-RAG is a promising basis for an intelligent, interpretable, and adaptive FLP algorithm recommendation method.
@@
This paper introduces a dynamic and domain-specific solution using the KG-RAG framework to generate data-driven algorithm recommendations to solve the FLP, given a user-defined query and a domain-specific knowledge graph of historical solutions. Our approach begins with constructing a formal KG that captures the complex relationships between FLP instances, their unique characteristics, and the performance of various solution methodologies. To leverage this structured knowledge base, the method employs a multi-faceted retrieval pipeline that integrates precise graph queries, flexible vector search, and high-level cluster analysis to gather comprehensive evidence. An LLM then synthesizes this evidence to generate data-driven and explainable recommendations for a complete methodology, including the algorithm, algorithm parameters, a suitable problem representation, and a constraint handling technique. This work includes an intelligent feedback loop, which allows the method to function as a continuously learning environment by incorporating new user-submitted solutions, thus ensuring its long-term relevance and accuracy.
@@
The rest of the paper has been structured as follows: Section 2 presents the proposed architecture and methodology, detailing the construction, domain-specific Knowledge Graph, and the multi-faceted retrieval and recommendation pipeline design. Section 3 provides a comparative experimental evaluation, benchmarking the KG-RAG method against a strong baseline LLM to empirically validate its performance on recommendation accuracy and its ability to generate data-driven, interpretable reasoning.
@@ 2 Architecture and methodology
The facility layout problem concerns placing a set of facilities (machines, departments, storage areas, etc.) inside a bounded site to optimise one or more performance measures. Practical FLP variations span multiple dimensions that materially affect solver choice and complexity. These begin with geometric representation, from the equal-area rectangles of early models [4] to the unequal area, aspect-ratio rectangles and general polygons of modern applications [1]. The problem scale is equally decisive [18], with the objective structure further constraining algorithm suitability, as single-objective methods, for example, minimizing material-handling cost, efficient area utilisation, minimizing material flow, may be ineffective for multi-objective trade-offs, for example, cost vs. safety vs. maintainability and vice versa [20]. Then the constraint coupling fundamentally shapes the feasibility landscape, including non-overlap, adjacency requirements, boundary constraints, and other geometric or operational limits. Together, these factors create a vast search space, leading to high variability in algorithm performance across different FLP instances. This diversity, as discussed in Section 2, underscores the necessity for a specialised algorithm recommendation method to adjust to each case’s unique features.
@@
Bridging this gap requires a systematic algorithm-selection approach that moves beyond generic recommendations. An ideal method must perform instance-aware matching, aligning suggestions with the specific scale, geometry, and constraints of the input account for objective type, ensuring alignment with single- or multi-objective goals and provide constraint-handling guidance, recommending methods such as repair operators, penalty functions, or exact feasibility checks. Beyond naming a method, it should supply validated algorithm parameters, such as population size for a genetic algorithm and starting temperature and cooling schedule for simulated annealing, as users often lack expertise in setting effective hyperparameters. For adoption, the method must ensure explainability and provenance, justifying recommendations with historical performance data, and deliver actionable outputs.
@@ 2.1 Architectural overview
A recommendation method is proposed to address the previously identified requirements. The method’s architecture, depicted in Figure 2, comprises several key components, each designed to address a specific research gap.
  • Initialisation of user query input: The process begins with the user defining their FLP. This component addresses the need for input flexibility, allowing users to specify diverse combinations of objectives, number of facilities, and constraints, moving beyond the rigid definitions.
  • Hybrid evidence retrieval: The method initiates a parallel, three-streamed retrieval process to overcome the limitations of static and single-faceted search methods. This component provides a comprehensive and adaptable search by collecting evidence from the Neo4j Knowledge Graph. This includes:
@@
  • Continuous learning via feedback: The design has a feedback loop. When users add new solved problem cases, the method learns and incorporates this information smartly, enriching the KG. This makes the method continuously learn and improve, allowing it to make more accurate and suitable recommendations in the future.
@@
The method is implemented in Python, utilizing a modern technology stack composed of Neo4j for the graph database and vector index, Google Gemini as the LLM and embedding engine, LangChain for orchestrating the RAG pipeline, and Gradio for the interactive web user interface.
@@ 2.2 Knowledge graph construction
Column Type Example
problem_id string P_6, D_10
num_facilities integer 6, 10
floor_W, floor_H float 30, 90
Problem representation string continuous space
facility_dimension_data string fixed dim fixed area
objective string min material handling cost
constraints string non-overlapping
constraint_handling string shapely intersection
method string GA, PSO, TabuSearch
Model_parameters String pop_size=50;n_gen=200;
cost float 1147.781, 18446.919
time_sec float 185.0, 290.19
source(reference) String Knez, M. and Gajsek, B.,
@@ 2.2.1 Data corpus and schema
The initial knowledge base was derived from a CSV file, a comprehensive dataset aggregating solved FLP instances from academic literature and benchmarks. Each row represents a unique problem-solution pair, capturing critical attributes as detailed in Table 1.
@@ 2.2.2 Graph schema design
(ConsHandling) - [IS_TYPE_OF] $\rightarrow$ (ConsHandlingCluster)
@@ 2.3.1 Multi-faceted evidence retrieval
    • Flexible filtering: The query applies an or constraint between constraints and goals to prevent the search from narrowing down too much. Besides that, it also dynamically regulates the quantity of facilities. Instead of searching for an exact amount, the Cypher query searches for problems in a ± 25% range of the number given by the user (i.e., between 0.75 * n and 1.25 * n). This dynamic range significantly increases the possibility of retrieving meaningful historical information, even if no exact number matches the facility number in the knowledge graph.
@@
  • Cluster-based search: To further support the provision of high-level statistical context, the method analyses the list of problem IDs found by the graph-based search (before applying the final limit). It does not aggregate queries to find the top three most common patterns to solutions for the Scale cluster and objective cluster associated with the user query. This gives strong trend-based evidence as well as the instance-level results.
2.3.2 LLM compilation and generation
@@
3 Comparative evaluation
This section presents the experimental design and results of a comparative study conducted to evaluate the performance of our proposed KG-RAG method. The primary objective is empirically confirming that grounding an LLM in a well-structured, domain-specific Knowledge Graph leads to significantly accurate and interpretable recommendations than a baseline LLM approach using raw, structured data.
3.1 Evaluation methodology
The evaluation is designed as a direct comparative analysis between our proposed model and a strong baseline model, with five distinct test cases designed to cover a range of real-world FLP scenarios, from simple, single-objective problems to more complex multi-objective and ill-defined queries. Finally, Section 4 concludes by summarizing the paper’s primary contributions and key findings, discussing the method’s current limitations, and outlining promising directions for future research.
  • Baseline model: The baseline model consists of the same Google Gemini 1.5 Flash LLM as the proposed model, but instead of a KG, it is provided with the entire CSV dataset file as its knowledge source. This represents a powerful, traditional RAG method that relies on the LLM’s ability to find and reason with information from a flat, tabular text file.
  • Ground truth establishment: A ground truth recommendation was established for each test case to create a fair benchmark for accuracy. This was not a subjective choice but the result of a systematic, manual analysis of the source dataset, following a process that mirrors the logic of ground truth finding.

### v2

@@ Abstract
Selecting a solution algorithm for the Facility Layout Problem (FLP), an NP-hard optimization problem with multiobjective trade-off, is a complex task that requires deep expert knowledge. The performance of a given algorithm depends on the specific characteristics of the problem, such as the number of facilities, objectives, and constraints. This creates a need for a data-driven recommendation method to guide algorithm selection in automated design systems. This paper introduces a new recommendation method to make this expertise accessible, based on a Knowledge Graph-Based Retrieval-Augmented Generation (KG-RAG) framework. In this framework, a domain-specific knowledge graph (KG) is constructed from the literature. The method then employs a multifaceted retrieval mechanism to gather relevant evidence from this KG using three distinct approaches: precise graph-based search, flexible vector-based search, and cluster-based high-level search. The retrieved evidence is utilized by a Large Language Model (LLM) to generate algorithm recommendations based on data-driven reasoning. This KG-RAG framework is tested on a use case consisting of six problems comprising of complex multi-objective and multi-constraint FLP case. The results are compared with the Gemini 1.5 Flash chatbot. The results show that KG-RAG achieves an average reasoning score of 4.7 out of 5 compared to 3.3 for the baseline chatbot.
Keywords: Facility Layout Problem, Algorithm Selection, Knowledge Graph, Retrieval-Augmented Generation, Large Language Models, Recommender systems.
@@ 1 Introduction
The Facility Layout Problem (FLP) is an important part of industrial engineering and operations research [1]. Because the operating cost and service performance are directly influenced by effective layout design. The suitable placement of facilities reduces the total operating costs by 20% to 50% [2]. The computational complexity of FLP stems from its combined discrete and continuous nature. In practice, this complexity is magnified by multi-objective trade-offs. For example, minimize material handling costs and maximize safety. This multi-objective problem is coupled with constraints such as aspect-ratio, adjacency and circulation constraints, and dynamic considerations such as reconfigurations to accommodate product-mix changes, demand changes, or equipment changes. These considerations render many FLP variants NP-hard, so exact mathematical programming is computationally tractable only for small or specially structured instances [3]. This FLP is also dynamic in nature because of evolving real-world demand. Selecting the algorithm for the FLP is challenging and requires expert knowledge because of the complex nature of the FLP.
@@
The solution techniques have been evolving as the complexity of the FLP increases, as mentioned above. Initial solutions, such as quadratic assignment problem [4], mixed-integer programming, and graph-theoretic formulations [5] for equal-area FLP. These approaches guarantee optimal solutions; however, the exponential growth in computational time severely limits their practical use in industrial-scale problems. As the problem size grows, researchers shifted to heuristic approaches, such as construction and improvement heuristics. Here, the construction heuristics build layouts from the ground up, for example, an automated layout design program and computerized relationship layout planning [6], whereas the improvement heuristics build solutions incrementally, for example computerized relative allocation of facilities technique [6]. These techniques are likely to converge to suboptimal layouts. To enhance the ability of handling large and complex search spaces, various metaheuristic techniques, including particle swarm optimization, genetic algorithms, ant simulated annealing, colony optimization, and others, have been introduced [7]. These techniques increase the likelihood of avoiding local optima. Moreover, the hybrid metaheuristic approaches, such as biased random-key genetic algorithms with linear programming repair, have recently been developed as state-of-the-art solutions, with enhanced solution quality and scalability [8].
@@
Despite these developments, the situation in FLP today is a broad and scattered collection of niche algorithms. This is typical of the no free lunch theorems, which state that when performance is averaged across all possible optimization problems, no single algorithm is universally superior. For example, an algorithm that performs well in small-scale, static settings may be desired in large-scale, multi-objective, or dynamic conditions [9]. As a result, attention has turned to the meta-problem of algorithm selection, which involves selecting the algorithm, representation, and parameters best suited to an instance of information [21].
@@
The early attempts to address this meta-problem were static decision-support methods. For instance, a GenOpt-based [10] framework was developed to guide practitioners in selecting from a wide range of general optimization algorithms. Their approach, which utilized flowcharts and choice matrices, was designed to assist novices by matching algorithm characteristics to problem types and performance requirements. While this represents a valuable step towards systematizing algorithm selection, the framework was not specific to the unique complexities of the FLP. Furthermore, like other static methods, it is not dynamic and cannot adapt to new algorithms or shifting problem definitions without manual updates [10]. To improve flexibility, research employed data-driven approaches, namely meta-learning, that predict algorithm performance from problem meta-features, for example, size, constraints, and statistical properties [11]. SATzilla, for example, selects the fastest solver from a portfolio based on instance features and has demonstrated strong performance in combinatorial optimization [12]. This approach later evolved into AutoML, where algorithm selection is combined with hyperparameter tuning under the Combined Algorithm Selection and Hyperparameter (CASH) framework [13]. Although adaptive frameworks and AutoML have built significant capabilities, the emergence of Large Language Models (LLMs) such as OpenAI’s GPT series and Google’s Gemini introduces new possibilities [14]. Trained on large, general-purpose corpora, these models are well-suited for acquiring language, reasoning, and in-context learning; however, when applied to domain-specific problems such as FLP optimization they can produce hallucinations (plausible but incorrect responses) a risk that is exacerbated by their weak domain adaptation [15] [16].
@@
AI research has increasingly focused on grounding LLMs with structured and verifiable knowledge to reduce hallucinations and improve domain adaptation. Within this direction, two complementary paradigms have gained prominence: Retrieval-Augmented Generation (RAG) and Knowledge Graphs (KG). RAG retrieves relevant domain-specific information before generation, mitigating hallucinations and enabling integration of up-to-date content without retraining [17]. KG organise domain entities and relationships, enabling explainable reasoning and mitigating data sparsity or cold-start issues. The relational pathways in KG enhance accuracy and interpretability, which are critical for decision support in technical domains [18]. A new approach integrates these paradigms. A Knowledge Graph-Based Retrieval-Augmented Generation (KG-RAG) [19] model uses a structured KG as the retrieval source for an RAG. The KG offers an explainable, linked domain model, and RAG allows LLMs to flexibly query and blend this knowledge into transparent and evidence-based proposals. This integration is suitable for comprehensive domain knowledge and justifiable output for complex recommendation tasks. The literature indicates a convergence of a long-standing operations research problem with a contemporary AI opportunity. Algorithmic fragmentation in FLP, as implied by the no free lunch principle, makes the task of algorithm selection a critical bottleneck. Current machine learning-based selectors, though adaptive, tend to operate as black boxes. Concurrently, the KG-RAG paradigm addresses Large Language Model (LLM) limitations of hallucinations and poor performance in domain adaptation while allowing for structured and transparent reasoning.
@@
This paper introduces a dynamic and domain-specific solution using the KG-RAG framework to generate data-driven algorithm recommendations for the FLP, given a user-defined query and a domain-specific KG of historical solutions. Our approach begins with constructing a formal KG that captures the complex relationships between FLP instances, their unique characteristics, and the performance of various solution methodologies. To leverage this structured knowledge base, the method employs a multi-faceted retrieval pipeline that integrates precise graph queries, flexible vector search, and high-level cluster analysis to gather comprehensive evidence. An LLM then synthesizes this evidence to generate data-driven and explainable recommendations for a complete methodology, including the algorithm, algorithm parameters, a suitable problem representation, and a constraint handling technique. This work includes an intelligent feedback loop, which allows the method to function as a continuously learning environment by incorporating new user-submitted solutions, thus ensuring its long-term relevance and accuracy.
@@
The paper has been structured as follows: Section 2 presents the proposed architecture and methodology, detailing the construction, domain-specific KG, and the multi-faceted retrieval and recommendation pipeline design. Section 3 provides a comparative experimental evaluation, benchmarking the KG-RAG method against the Gemini 1.5 Flash LLM chatbot to empirically validate its performance on recommendation accuracy and its ability to generate data-driven, interpretable reasoning.
@@ 2 Architecture and methodology
The FLP concerns placing a set of facilities (machines, departments, storage areas, etc.) inside a bounded site to optimize one or more performance measures. While Figure 1 provides a simple illustration of the FLP layout context, the practical problem spans multiple dimensions that materially affect solver choice and complexity. These begin with geometric representation, from the equal-area rectangles of early models [4] to the unequal area, aspect-ratio rectangles and general polygons of modern applications [1]. The problem scale is equally decisive [18], with the objective structure further constraining algorithm suitability, as single-objective methods, for example, minimizing material-handling cost, efficient area utilisation, minimizing material flow, may be ineffective for multi-objective trade-offs, for example, cost vs. safety vs. maintainability and vice versa [20]. Then the constraint coupling fundamentally shapes the feasibility landscape, including non-overlap, adjacency requirements, boundary constraints, and other geometric or operational limits. Together, these factors create a vast search space, leading to high variability in algorithm performance across different FLP instances. This diversity, as discussed in Section 1, underscores the necessity for a specialised algorithm recommendation method to adjust to each case’s unique features.
@@
Bridging this gap requires a systematic algorithm-selection approach that moves beyond generic recommendations. An ideal method must perform instance-aware matching, aligning suggestions with the specific scale, geometry, and constraints of the input account for objective type, ensuring alignment with single- or multi-objective goals and providing constraint-handling guidance, recommending methods, for example, repair operators, penalty functions, or exact feasibility checks. Beyond naming a method, it should supply validated algorithm parameters, such as population size for a genetic algorithm and starting temperature and cooling schedule for simulated annealing, as users often lack expertise in setting effective hyperparameters. For adoption, the method must ensure explainability and provenance, justifying recommendations with historical performance data, and deliver actionable outputs.
@@ 2.1 Architectural overview
A recommendation method is proposed to address the previously identified requirements. The method’s architecture, depicted in Figure 2, comprises several key components, each designed to address a specific research gap, as explained below.
  • Initialisation of user query input: The process begins with the user defining their FLP. This component addresses the need for input flexibility, allowing users to specify diverse combinations of objectives, the number of facilities, and constraints, moving beyond the rigid definitions.
  • Hybrid evidence retrieval: The method initiates a parallel, three-streamed retrieval process to overcome the limitations of static and single-faceted search methods. This component provides a comprehensive and adaptable search by collecting evidence from the Neo4j KG. This includes:
@@
  • Continuous learning via feedback: The design has a feedback loop. When users add new solved problem cases, the method learns and incorporates this information smartly, enriching the KG.
@@
The method is implemented in Python, utilizing a modern technology stack composed of Neo4j for the graph database and vector index, Google Gemini as the LLM and embedding engine, LangChain for orchestrating the RAG pipeline, and Gradio for the interactive web user interface. The user interacts with the system through the interface shown in Figure 3. They can input specific problem parameters, such as the number of facilities and key constraints, to receive a comprehensive methodology recommendation, including the selected algorithm, its parameters, and the underlying data-driven reasoning.
Figure 3: User Interface
@@ 2.2 Knowledge graph construction
Column Description Type
Problem ID Problem identifier string
Number of Facilities Number of Facilities integer
Floor Dimensions Width × Height (m) float
Problem Representation Layout type string
Facility Data Dimensions type string
Objective Optimization goal string
Constraints Layout restrictions string
Constraint Handling How constraints are handled string
Optimization Method Solver used string
Model Parameters Key algorithm parameters String
Solution Cost Solution cost float
Computation Time Computation time float
Source Reference String
@@ 2.2.1 Data corpus and schema
The initial knowledge base was derived from a CSV file, a comprehensive dataset aggregating solved FLP instances from academic literature and benchmarks. Table 1 summarizes the structure of this dataset, where each row represents a layout problem and its solution. The columns capture key information such as the problem ID, number of facilities, floor dimensions, problem type, constraints, solver method, algorithm parameters, solution cost, computation time, and source reference.
@@ 2.2.2 Graph schema design
(ConsHandling) - [is_type_of] $\rightarrow$ (ConsHandl Cluster)
@@ 2.3.1 Multi-faceted evidence retrieval
    • Flexible filtering: The query applies an or constraint between constraints and goals to prevent the search from narrowing down too much. Besides that, it also dynamically regulates the quantity of facilities. Instead of searching for an exact amount, the Cypher query searches for problems in a ± 25% range of the number of facilities given by the user. This dynamic range significantly increases the possibility of retrieving meaningful historical information, even if no exact number matches the facility number in the KG.
@@
  • Cluster-based search: To further support the provision of high-level statistical context, the method analyses the list of problem IDs found by the graph-based search (before applying the final limit). It then aggregates queries to find the top three most common patterns in solutions for the scale cluster and objective cluster associated with the user query. This provides both strong trend-based evidence and instance-level results.
2.4 LLM compilation and generation
@@
2.5 Comparative evaluation
This section presents the experimental design and results of a comparative study conducted to evaluate the performance of our proposed KG-RAG method. The primary objective is empirically confirming that grounding an LLM in a well-structured, domain-specific KG leads to significantly accurate and interpretable recommendations than a baseline LLM approach using raw, structured data.
2.5.1 Evaluation methodology
The evaluation is designed as a direct comparative analysis between our proposed model and a strong baseline model, with five distinct test cases designed to cover a range of real-world FLP scenarios, from simple, single-objective problems to more complex multi-objective and ill-defined queries.
  • Baseline model: The baseline Google Gemini 1.5 Flash LLM as the proposed model, but instead of a KG, it is provided with the entire CSV dataset file as its knowledge source. This represents a powerful, traditional RAG method that relies on the LLM’s ability to find and reason with information from a flat, tabular text file.
  • Ground truth establishment: For each test case, a ground truth recommendation was established to provide a fair benchmark for accuracy. This was not a subjective choice but the outcome of a systematic manual analysis of the source dataset. Figure 4 illustrates the process, which mirrors the logic of ground truth determination.